# Neuromorphic force-control in an industrial task: validating energy and latency benefits

### v1

@@ Abstract
As robots become smarter and more ubiquitous, optimizing the power consumption of intelligent compute becomes imperative towards ensuring the sustainability of technological advancements. Neuromorphic computing hardware makes use of biologically inspired neural architectures to achieve energy and latency improvements compared to conventional von Neumann computing architecture. Applying these benefits to robots has been demonstrated in several works in the field of neurorobotics, typically on relatively simple control tasks. Here, we introduce an example of neuromorphic computing applied to the real-world industrial task of object insertion. We trained a spiking neural network (SNN) to perform force-torque feedback control using a reinforcement learning approach in simulation. We then ported the SNN to the Intel neuromorphic research chip Loihi interfaced with a KUKA robotic arm. At inference time we show latency competitive with current CPU/GPU architectures, two orders of magnitude less energy usage in comparison to traditional low-energy edge-hardware. We offer this example as a proof of concept implementation of a neuromoprhic controller in real-world robotic setting, highlighting the benefits of neuromorphic hardware for the development of intelligent controllers for robots.
@@ I INTRODUCTION
Moving towards the integration of intelligent robots in daily life, sustainability considerations require stricter optimization of energy consumption not only for actuation, but also for the computing involved in robot control. The current energy requirements of GPUs and CPUs either limit the scale and “intelligence” of edge computation, or require high-latency communication protocols for cloud computing. While moving an AI system such as IBM Watson or ChatGPT to the edge requires killowatts of power for continuous inference111Ignoring training energy costs. [1, 2], the human brain accomplishes much more complex behaviour at a tiny fraction of this [3]. Biological inspiration therefore continues to influence both algorithm and hardware development in the field of neuromorohic computing, with an increasing number of applications in robotics.
@@
To address this, we present a real-world neurorobotic system in which a robotic arm equipped with a force-torque (FT) sensor accomplishes an insertion task while fully controlled by neuromorphic hardware. The classic robotic insertion task has been chosen as it serves as the foundation for a number of related sub-tasks such as screw insertion, cable attachment, part assembly, etc. Additionally, the “peg-in-hole” task has historically served as a benchmark for robotic integration of new algorithms and hardware [30, 31, 32, 33, 34, 35, 36], etc (see [37] for an overview).
@@
This RL in simulation + sim2real technique is similar to that recently applied to drone and quadrupedal maneuvering using conventional ANNs [38, 39, 40, 41]. Our approach is similar, but realised with SNNs and deployed on a neuromorphic chip. We argue that, as the hardware matures, this could lead to energy efficient trained controllers for a broad deployment of intelligent robots. Our experiments have shown two orders of magnitude less energy usage in comparison to current SotA non-neuromorphic edge-optimized hardware at a similar latency.
We highlight the following novel contributions of this work: this is the first real-world non-trivial robotic use-case (peg-in-hole insertion with an industrial robotic arm) fully guided by neuromorphic hardware. In this setting, energy and latency measurements of neuromorphic hardware were herein performed for the first time. We have found energy spent on running the network on the neuromorphic chip to be in the milliwatt range. Moreover, this is the first neurorobotic system to incorporate a force-torque sensor.
@@ II Methodology
We first develop a simulation of the robotics setup, train the neural network based controller with spiking-RL using Pytorch, port the network to neuromorphic hardware, and finally run the trained policy on the real robot applying various techniques to address the Sim2Real gap. A visualization of these steps is shown in Figure 1.
@@ II-A Simulation Setup
The setup is simulated within the Neurorobotics Platform (NRP) [16], developed as part of the Human Brain Project [42] and based on ROS [43] and Gazebo [44]. A robotic arm with 7 DOF mounted to a table is modelled. The end-effector consists of a cylindrical peg with a flat end and a 6-axis FT sensor. The table consists of a board with a circular hole. The peg can fit in the hole with a clearance of . This setup closely resembles the one used by previous teams (i.e. [34]). The setup is shown in Figure 2. The full details of the simulation setup and underlying programming have been outlined in a separate complimentary publication [45]. It primarily explores the implementation of RL within the NRP.
@@ II-D1 Controller
A low level compliant controller was designed and implemented both for simulation as well as for the real robot555See footnote 2. More precisely, a Cartesian Impedance Controller (CIC) was implemented and tuned using the methodology outlined in [49]. The controller was implemented with the KUKA Fast Research Interface (FRI).
@@ III-B Real-World Performance
When the policy is initially run on the real robot (Figure 1, right), we find an insertion success rate of 0% (across 10 runs, where the system is deemed to have failed if it does not insert within 30 s). After the Sim2Real techniques (Section II-D3) we find a 100% insertion success rate across 50 runs. Using success rate to compare the simulated and real implementations is therefore not possible, so we instead use the time to insertion (shown in Table III-B). All values are calculated across 50 runs, and the peg starts its search roughly 2 cm from the hole (in a random direction).
A similar median and minimum insertion time is observed. Note that the mean real insertion time is roughly double the simulated time. This is due to a small number of outlier runs in which the peg has trouble navigating the complex real friction profile and temporarily gets stuck (which also explains the maximum insertion time discrepancy). Yet, perfect friction modelling is not our goal, and the a few second median time is comparable with SotA implementations (see, for example, Figure 16 in [36]). We now move to the energy and latency.
Simulation |
Real |
Mean |
3.7 s |
8.4 s |
Median |
3.4 s |
5.3 s |
Minimum |
2.5 s |
2.2 s |
Maximum |
7.4 s |
28.1 s |
@@ III-C Latency & Energy
Hardware |
Latency [] |
CPU |
GPU |
Loihi 1* |
@@
CPU: 11th Gen Intel® Core™ i7-1165G7 @ 2.80GHz × 8.
GPU: NVIDIA®/ Mesa Intel® Xe Graphics (TGL GT2) / GeForce MX550
* Note: Some values are not available on Loihi 1 due to probe limitations.
@@ IV-A Latency & Energy
Contrasting these milliwatt-range power values with SotA non-neuromorphic edge computational hardware, the potential benefits are evident. Currently, optimized edge-hardware requires a few watts of power for similar tasks [53, 54]. Neuromorphic hardware could therefore allow for more complex computations on autonomous robots, as well as orders-of-magnitude longer times between recharging. Additionally, non-neuromorphic edge-hardware often achieves low power consumption at the cost of latency (generally on account of cloud communication). For instance, [53] cite latencies on the order of hundreds of milliseconds, compared to a few milliseconds by computing directly on Loihi.
@@
The measured energy and latency results are comparable to those reported in simulated neuromorphic use cases, hence validating the idea that the benefits of neurobotic systems continue to hold when moved from simulations and simplistic use cases to more complex real-world applications. For example, [19] reports 200 and 1 per inference using a slightly larger network on Loihi 1 for simulated robotic control. [12] run a 4-layer SNN on Loihi 1D drone control. Although energy values are not reported, they do find a latency of 0.05 per step, corresponding to an equivalent of 0.5 per inference. This value may be smaller than what we find on account of especially sparse data (coming from a neuromorphic camera). We noted that the latency values for CPU and GPU vary largely between devices. Indeed, [19] cite 15 latency on CPU, 10 times larger than our value. Ultimately we find similar latency between CPU and neuromorphic hardware.
Previous RL implementations of peg-in-hole (e.g. [34], [55], [31]) generally invoke ANNs of a similar size to the SNN introduced here. Therefore, we should expect that the inference-time energy improvements over CPU and GPU (Table III-C) apply here. Of course, inference represents only one part of the entire robotic control system. The true benefits of neurorobotics will only be unlocked when the robots themselves — as with our bodies — also operate on spiking architectures. This would avoid losing energy and latency to encoding and decoding spikes between the neuromorphic hardware and non-neuromorphic robot controller.
@@
It should be noted that Table III-C refers to the computation times on-chip. On account of a known I/O bottleneck in Loihi 1 the robot-chip communication, which in theory should be negligible, ultimately increases the latency to 16 ms. Although addressed in Loihi 2 [56], we do not have an on-site physical chip, and therefore cannot measure the end-to-end latency on account of the cloud-communication cost. There is no reason, however, that the end-to-end latencies should be larger than what is quoted in Table III-C.
IV-B Learned Movement
@@
The insertion times (Table III-B) are comparable between simulation and the real robot. We find quantization effects on Loihi (previously explored by [57]) to be negligible. We notice that, even when applying domain randomization in simulation training to vary the friction coefficient (section II-D3) the movement behaviour still differs somewhat due to the complex friction profile between the peg and surface. Notably, on the real surface, the peg will often initially not move, and then jerk forward when the friction is overcome. This jerking was not captured in simulation, although insertion on the real robot is still successful regardless.
IV-C Application Specific / Peg-in-hole
@@
We had initially hypothesized that the arm, using the FT sensor and ability to move in angular space, would learn to feel out the force-torque profile corresponding to being on the edge of the hole. We find that the trained policy is such that the angular movement is quite small (° for all three axes). This means that, generally the peg is held approximately vertical. Additionally, during the exploration phase, the trained policy returns actions sending the end-effector back-and-forth into the positive and negative and directions (relative to the hole). Although the peg appears to sometimes react to arriving at the edge of the hole (by moving towards the hole centre), in a number of runs we certainly see the peg slipping into the hole, likely by accident, from this learned sweeping behaviour. However, in the case where the peg only slips part-way into the hole, the slight angular movements do appear to help it arrive at the bottom.
@@
IV-D Neurorobotic Outlook
We have demonstrated robotic insertion using a neuromorphic high-level controller as a proof of concept, meant to emphasize that neuromorphic computing is no longer simply up-and-coming, but rather sufficiently mature to move towards tackling real robotic use cases.
The energy benefits are evident, and although we find latency to be similar to classical devices, it is important to remember that Loihi is intended as a flexible research platform and has not been optimized for specific tasks. Indeed, such platforms can already be used as test-beds and prototyping tools for current non-research neuromorphic and mixed-signal devices with SoTA performance666eg. DYNAP-SE, ROLLS, Innaterra, etc.. Neuromorphic hardware also allows for “online” and “continuous” learning unique to SNNs and inspired by the human brain which, although found not to be needed for successful insertion in our use case, we nonetheless hope to explore in a future publication.
This implementation is also intended to act as a software base for future neurorobotic development. As with the first RL peg-in-hole implementation, we also hope that this first neuromorphic implementation contributes to an exciting new domain of robotics and automatization — that of neurorobotics.

### v3

@@ Abstract
As robots become smarter and more ubiquitous, optimizing the power consumption of intelligent compute becomes imperative towards ensuring the sustainability of technological advancements. Neuromorphic computing hardware makes use of biologically inspired neural architectures to achieve energy and latency improvements compared to conventional von Neumann computing architecture. Applying these benefits to robots has been demonstrated in several works in the field of neurorobotics, typically on relatively simple control tasks. Here, we introduce an example of neuromorphic computing applied to the real-world industrial task of object insertion. We trained a spiking neural network (SNN) to perform force-torque feedback control using a reinforcement learning approach in simulation. We then ported the SNN to the Intel neuromorphic research chip Loihi interfaced with a KUKA robotic arm. At inference time we show latency competitive with current CPU/GPU architectures, and one order of magnitude less energy usage in comparison to state-of-the-art low-energy edge-hardware. We offer this example as a proof of concept implementation of a neuromoprhic controller in real-world robotic setting, highlighting the benefits of neuromorphic hardware for the development of intelligent controllers for robots.
@@ I INTRODUCTION
Moving towards the integration of intelligent robots in daily life, sustainability considerations require stricter optimization of energy consumption not only for actuation, but also for the computing involved in robot control. The current energy requirements of GPUs and CPUs either limit the scale and “intelligence” of edge computation, or require high-latency communication protocols for cloud computing. While moving an AI system such as IBM Watson or ChatGPT to the edge requires killowatts of power for continuous inference111Ignoring training energy costs. [1, 2], the human brain accomplishes much more complex behaviour at a tiny fraction of this [3]. Biological inspiration therefore continues to influence both algorithm and hardware development in the field of neuromorphic computing, with an increasing number of applications in robotics.
@@
To address this, we present a real-world neurorobotic system in which a robotic arm equipped with a force-torque (FT) sensor accomplishes an insertion task while fully controlled by neuromorphic hardware. The classic robotic insertion task has been chosen as it serves as the foundation for a number of related sub-tasks such as screw insertion, cable attachment, part assembly, etc. Additionally, the “peg-in-hole” task has historically served as a benchmark for robotic integration of new algorithms, sensors, and hardware [30, 31, 32, 33, 34, 35, 36], etc (see [37] for an overview).
@@
This RL in simulation + sim2real technique is similar to that recently applied to drone and quadrupedal maneuvering using conventional ANNs [38, 39, 40, 41]. Our approach is similar, but realised with SNNs and deployed on a neuromorphic chip. We argue that, as the hardware matures, this could lead to energy efficient trained controllers for a broad deployment of intelligent robots. Our experiments have shown one order of magnitude less energy usage in comparison to current SotA non-neuromorphic edge-optimized hardware at a similar latency.
We highlight the following novel contributions of this work: this is the first real-world non-trivial robotic use-case (peg-in-hole insertion with an industrial robotic arm) fully guided by neuromorphic hardware. In this setting, energy and latency measurements of neuromorphic hardware were herein performed for the first time. We have found energy spent on running the network on the neuromorphic chip to be in the microjoule range. Moreover, this is the first neurorobotic system to incorporate a force-torque sensor.
@@ II Methodology
We first develop a simulation of the robotics setup, train the neural network based controller with spiking-RL using Pytorch, port the network to neuromorphic hardware, and finally run the trained policy on the real robot, applying various techniques to address the Sim2Real gap. A visualization of these steps is shown in Figure 1.
@@ II-A Simulation Setup
The setup is simulated within the Neurorobotics Platform (NRP) [16], developed as part of the Human Brain Project [42] and based on ROS [43] and Gazebo [44]. A robotic arm with 7 DOF mounted to a table is modelled. The end-effector consists of a cylindrical peg with a flat end and a 6-axis FT sensor. The table consists of a board with a circular hole. The peg can fit in the hole with a clearance of , and is initialized as per a Gaussian distribution centered on the hole with . This setup closely resembles the one used by previous teams (i.e. [34]). The setup is shown in Figure 2. The full details of the simulation setup and underlying programming have been outlined in a separate complimentary publication [45]. It primarily explores the implementation of RL within the NRP.
@@ II-D1 Controller
A low level compliant controller was designed and implemented both for simulation as well as for the real robot. More precisely, a Cartesian Impedance Controller (CIC) was implemented and tuned using the methodology outlined in [49]. The controller was implemented with the KUKA Fast Research Interface (FRI).
@@ III-B Real-World Performance
When the policy is initially run on the real robot (Figure 1, right), we find an insertion success rate of 0% (across 10 runs, where the system is deemed to have failed if it does not insert within 30 s). After the Sim2Real techniques (Section II-D3) we find a 100% insertion success rate across 50 runs. Using success rate to compare the simulated and real implementations is therefore not possible, so we instead use the time to insertion (shown in Table III-B). All values are calculated across 50 runs.
A similar median and minimum insertion time is observed. Note that the mean real insertion time is roughly double the simulated time. This is due to a small number of outlier runs in which the peg has trouble navigating the complex real friction profile and temporarily gets stuck (which also explains the maximum insertion time discrepancy). Yet, perfect friction modelling is not our goal, and median insertion time on the order of a few seconds is comparable with SotA implementations (see, for example, Figure 16 in [36]). We now move to the energy and latency.
Simulation | Real | |
Mean | 3.7 s | 8.4 s |
Median | 3.4 s | 5.3 s |
Minimum | 2.5 s | 2.2 s |
Maximum | 7.4 s | 28.1 s |
@@ III-C Latency & Energy
Hardware | [] | Latency [] |
CPU † | ||
GPU | a few 100‡ | * |
Loihi 1** | — | |
@@
† CPU: 11th Gen Intel® Core™ i7-1165G7 @ 2.80GHz × 8.
‡ NVIDIA® / Pascal 256 CUDA cores @ 1300 MHz / Jetson TX2 ([53] Alg. 1 + personal communication with author)
* NVIDIA® / Mesa Intel® Xe Graphics (TGL GT2) / GeForce MX550
** Note: Energy values are not available on Loihi 1 due to probe limitations.
@@ IV-A Latency & Energy
Contrasting the recorded Loihi energy values with SotA non-neuromorphic edge computational hardware (Table III-C), the potential benefits are evident. Currently, optimized edge-hardware requires energy on the order of hundreds of per inference for similar tasks/networks [53, 54, 55], an order of magnitude more than neuromorphic hardware. Such hardware could therefore allow for more complex computations on autonomous robots, as well as orders-of-magnitude longer times between recharging. Additionally, non-neuromorphic edge-hardware often achieves low power consumption at the cost of latency (generally on account of cloud communication). For instance, [54] cite latencies on the order of hundreds of milliseconds, compared to a few milliseconds by computing directly on Loihi (Figure 6). It is also important to remember that Loihi is intended as a flexible research platform and has not been optimized for specific tasks.
Previous RL implementations of peg-in-hole (e.g. [34, 56, 31]) generally invoke ANNs of a similar size to the SNN introduced here. Therefore, we should expect that the inference-time energy improvements over CPU and GPU (Table III-C) apply here. Of course, inference (and computation in general) represents only one part of the entire robotic control system. For larger robots (10s of kg), the power requirements of the mechanical actuators are roughly one order of magnitude larger than the power requirements of computation [57]. Yet for sub-kilogram robots computation is likely to be a non-negligible source of power-consumption. As edge-robots become smaller and the tasks we expect them to perform become more complex, efficient compute will become even more imperative.
@@
The measured energy and latency results are comparable to those reported in simulated neuromorphic use cases, hence validating the idea that the benefits of neurorobotic systems continue to hold when moved from simulations and simplistic use cases to more complex real-world applications. For example, [19] reports 200 and 1 per inference using a slightly larger network on Loihi 1 for simulated robotic control. [12] run a 4-layer SNN on Loihi 1D drone control. Although energy values are not reported, they do find a latency of 0.05 per step, corresponding to an equivalent of 0.5 per inference. This value may be smaller than what we find on account of especially sparse data (coming from a neuromorphic camera). We noted that the latency values for CPU and GPU vary largely between devices. Indeed, [19] cite 15 latency on CPU, 10 times larger than our value. Ultimately we find similar latency between CPU and neuromorphic hardware.
@@
It should be noted that Table III-C refers to the computation times on-chip. On account of a known I/O bottleneck in Loihi 1 the robot-chip communication, which in theory should be negligible, ultimately increases the latency to 16 ms. Although addressed in Loihi 2 [58], we do not have an on-site physical chip, and therefore cannot measure the end-to-end latency on account of the cloud-communication cost. There is no reason, however, that the end-to-end latencies should be larger than what is quoted in Table III-C. We also remind the reader that the energy per inference represents only one part of the overall energy required for computation. Loihi currently requires significant power for overhead operations (such as for spike I/O - handled by a separate FPGA). These are not included, as they are expected to be optimized with subsequent chip releases, to the point that the inference energy is primary (as was the case for classical chips).
IV-B Application Specific / Peg-in-hole
@@
The insertion times (Table III-B) are comparable between simulation and the real robot. We find quantization effects on Loihi (previously explored by [59]) to be negligible. We notice that, even when applying domain randomization in simulation training to vary the friction coefficient (Section II-D3) the movement behaviour still differs somewhat due to the complex friction profile between the peg and surface. Notably, on the real surface, the peg will often initially not move, and then jerk forward when the friction is overcome. This jerking was not captured in simulation, although insertion on the real robot is still successful regardless. Velocity and acceleration limits are the standard KUKA preset values. We do not expect varying the limits would affect the outcome, although this would need to be explicitly tested.
@@
We had initially hypothesized that the arm, using the FT sensor and ability to move in angular space, would learn to feel the force-torque profile corresponding to being on the edge of the hole. We find that the trained policy is such that the angular movement is quite small (° for all three axes). This means that generally the peg is held approximately vertical. Additionally, during the exploration phase, the trained policy returns actions sending the end-effector back-and-forth into the positive and negative and directions (relative to the hole). Although the peg appears to sometimes react to arriving at the edge of the hole (by moving towards the hole centre), in a number of runs we certainly see the peg slipping into the hole, likely by accident, from this learned sweeping behaviour. However, in the case where the peg only slips part-way into the hole, the slight angular movements do appear to help it arrive at the bottom.
@@
IV-C Neurorobotic Outlook
We have demonstrated robotic insertion using a neuromorphic high-level controller as a proof of concept, meant to emphasize that neuromorphic computing is no longer simply up-and-coming, but rather sufficiently mature to move towards tackling real robotic use cases. This implementation is additionally intended to act as a software base for future neurorobotic development.
Indeed, flexible research platforms such as Loihi can already be used as test-beds and prototyping tools for current non-research neuromorphic and mixed-signal devices with SoTA performance555eg. DYNAP-SE, ROLLS, Innaterra, etc.. Neuromorphic hardware also allows for “online” and “continuous” learning unique to SNNs and inspired by the human brain which, although found not to be needed for successful insertion in our use case, we nonetheless hope to explore in a future publication.
Looking forward, it is worth mentioning that current actuators do not operate on spiking principles. With neuromorphic hardware and sensors already in existence, spike-based actuators would allow for neurorobots in which energy and latency is not lost to encoding and decoding spikes between neuromorphic and non-neuromorphic components. Additional research is here needed to unlock the true potential of neurorobots.
Ultimately, and as with the first RL peg-in-hole implementation, we hope that this first neuromorphic implementation contributes to an exciting new domain of robotics and automatization — that of neurorobotics.